0 / 0 = ?

0 / 0 = ?

Grade:12

2 Answers

Ankit Sharma
11 years ago

0/0=undefined

0/R=0

R/0=Infinity

Hope this solves your problem

Revanth - -
11 years ago

IT IS AN INDETERMINATE FORM

 

O/O IS INDETERMINATE FORM 

LET O/O IS SOME X

NOW 0=(X)0 NOW THIS EQUATION IS CORRECT FOR MANY VALUES OF X SO THIS FORM IS INDETERMINATE U WILL STUDY THIS IN UR LIMITS CHAPTER

APPROVE IT IT IS CORRECT